OpenStudy (anonymous):

find the limit of sqrt. of (x^2+9) -5/ x+4 as x approaches -4

OpenStudy (anonymous):

\[\lim_{x\to-4}\frac{\sqrt{x^2+9}-5}{x+4}~~?\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

\[\lim_{x\to-4}\frac{\sqrt{x^2+9}-5}{x+4}\cdot\frac{\sqrt{x^2+9}+5}{\sqrt{x^2+9}+5}\] \[\lim_{x\to-4}\frac{x^2+9-25}{(x+4)\left(\sqrt{x^2+9}+5\right)}\] \[\lim_{x\to-4}\frac{x^2-16}{(x+4)\left(\sqrt{x^2+9}+5\right)}\] \[\lim_{x\to-4}\frac{(x+4)(x-4) }{(x+4)\left(\sqrt{x^2+9}+5\right)}\]

OpenStudy (anonymous):

you gotta be kidding me.... thats what i did and my roommate said i couldnt do it that way

OpenStudy (anonymous):

Tell him/her you're better than him/her.

OpenStudy (anonymous):

haha thanks

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

You're welcome!